OpenStudy (bbb911):

This is Geometry: In an isosceles triangle, the base is 8 inches long and the legs are 10 inches long. Find the measure of a base angle.

OpenStudy (anonymous):

the perpendicular to the base cuts the base in two. so cosine of the base angle is 1/2 * 8 / 10 = 2/5 so cos base angle = 0.4 - a job for ur calculator

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

you need trigonometry to solve this as well as geometry

OpenStudy (bbb911):

here is a multiple choice question A. 84 degrees B. 66 degrees C. 10 degrees D. 6 degrees

OpenStudy (anonymous):

thats the possible answers right? but whats the question?

OpenStudy (bbb911):

the question right up above

OpenStudy (anonymous):

oh yes ! sorry - but haven't you got a calculator to do this bbb? It will be good practice for you.

OpenStudy (anonymous):

have u understood the method i've used?

OpenStudy (bbb911):

umm do i have to use a calculator? because i dont have a scientific one :o

OpenStudy (anonymous):

u need a scientific one to do this u need to find the angle whose cosine is 0.4 the answer is 66 degrees

OpenStudy (bbb911):

oh yes i see now yea needs me to buy one thank you for your help i will be your fan :D

OpenStudy (anonymous):

ty - but buy one - they are very cheap

OpenStudy (bbb911):

oh ok thanks will do :)